OpenStudy (anonymous):

Given a triangle with b = 7 c = 3 and A = 37 degrees, what is the length of a?

OpenStudy (anonymous):

sounds like a job for the law of cosines... http://www.mathsisfun.com/algebra/trig-cosine-law.html

OpenStudy (anonymous):

I did that, but I keep getting an answer that isn't an answer choice

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

6.5

OpenStudy (anonymous):

okay, just a sec...

OpenStudy (anonymous):

 \[a^2=b^2+c^2-2bc \cos A= 7^2+3^2-2(7)(3)\cos 37^{\circ}=49+9-42(.7986355...)\] \[=24.457308...\Rightarrow a=\sqrt{a^2}=\sqrt{24.457308...}=4.945433103...\]

OpenStudy (anonymous):

did you find your mistake?

OpenStudy (anonymous):

Yeah I did. Thank you so much

OpenStudy (anonymous):

you're welcome!